# The Euler class of the normal bundle of a Seifert fibration and horizontal foliations

### v1

@@ 2.1. Seifert notation
Consider $M$ constructed by attaching $n+1$ solid tori $N_{i}$, $i=0,1,\dots,n$, to a trivial $S^{1}$-bundle over an orientable surface $F$ with $|\partial F|=n+1$, where we assume that $N_{0}$ is a tubular neighbourhood of a regular fibre and the $N_{i}$ are tubular neighbourhoods of the singular fibres for $i=1,2,\ldots,n$. Denote this trivial $S^{1}$-bundle by $P\cong F\times S^{1}$. Fix orientations on $F$ and on the fibres so that they intersect positively.
@@
It follows from [8, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle. Hence $\mbox{kernel}(f^{*})$ is generated by $bu$. Then by (2.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $b$, which is condition (2.1.1).
@@
    From $\sigma$ one can define an obstruction $2$-cochain, denoted by $c_{\sigma}$, that represents the relative Euler class of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$. By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$. (See [9, Section 2] for instance.)
@@ 3.1. The second cohomology group of $\Gamma$
Jankins and Neuman determined a canonical isomorphism between $H^{2}(\Gamma)$ and the abelian group having generators $A_{0},A_{1},\ldots,A_{n}$ and relations $A_{0}=b_{i}A_{i}$ for $\leq i\leq n [11$]. (The case that $g=0$ is discussed in [11] and the general case in [10].) Each element of $H^{2}(\Gamma)$ can be uniquely expressed in the form
@@ 3.2. The normal bundle of the Seifert fibring of $M$ and discrete representations
$\psi(\gamma)(x,y)=(\gamma\cdot x,\varphi(\gamma)(y))$
@@
Now proceed as in the proof of [2, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [2, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$. Thus
@@
It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [14, Lemma 2].
@@
We saw in (3.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$. Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [14, Lemma 2]. It follows that $\pi_{1}(M)$ is left-orderable [6, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [6, Theorem 1.3]. ∎
4. Examples and applications
@@
That any co-oriented taut foliation on $M$ must be horizontal is [5, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general. Thus (1) holds. Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2). ∎
@@
Example 4.2 (Dehn surgeries on the trefoil knot).
Let $K=T(2,3)$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$. There is a natural Seifert fibre structure on the exterior $X=S^{3}\setminus\nu(K)$, whose regular fibre has slope $6\mu+\lambda$ on $\partial X$. Hence, $p/q$-surgery on $K$ yields a small Seifert fibred space whenever $p/q\neq 6$.
It is known that $X(p/q)$ admits co-oriented taut foliations when $p/q<1$ [16, §3]. On the other hand, [9, Theorem 1.7] shows that the Euler class of the co-oriented taut foliation constructed by Roberts in [16] vanishes if and only if
@@
Combining Example 4.2 with Lemma 4.1 yields Theorem 1.6.
@@
If $e(\nu_{M})=0$, then $d=\gcd(b_{i},b_{j})\implies a_{i}\equiv a_{j}\mbox{ {\rm(mod $d$)}}$.
@@
if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 2.1(1). Equivalent sets of Seifert invariants yield homeomorphic manifolds.
@@ 4.3. $\mathcal{B}$ is hyperbolic
4.4. Small values of $|H_{1}(M)|$
It is known that if $|H_{1}(M)|\leq 2$, then the Euler class $e(\nu_{M})$ vanishes (see [9, Proposition 2.1] for instance). In this section, we show that this fact can also be obtained from the computations in this article.
We have already discussed the case $M=M(0;b)$ in Remark 2.2, so we assume that $n>0$.
We recall the presentation of $H^{2}(\pi_{1}(\mathcal{B}))$ in §3.1, where $H^{2}(\pi_{1}(\mathcal{B}))$ is generated by $A_{0},A_{1},\ldots,A_{n}$ with relations $A_{0}=b_{i}A_{i}$ for $i=1,\ldots,n$. In §3, we showed that $e(\nu_{M})=\alpha^{*}(e(\overline{\varphi}))=0$, where $\overline{\varphi}\colon\pi_{1}(\mathcal{B})\to PSL(2,\mathbb{R})$ is discrete and faithful, if and only if $\exists\,m\in\mathbb{Z}$ such that $e(\overline{\varphi})=m\,e(E_{M})$. See (3.1.2) and (3.1.3). When $M=M\bigl(0;\,b,\tfrac{a_{1}}{b_{1}},\ldots,\tfrac{a_{n}}{b_{n}}\bigr)$, by (3.1.1) and (3.2.2),
$e(E_{M})=-\bigl(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n}\bigr),$
and
$e(\overline{\varphi})=-\left((2-n)A_{0}+A_{1}+A_{2}+\cdots+A_{n}\right)$
Suppose that $m$ is an integer such that $ma_{i}\equiv 1\,(\text{{\rm mod} }b_{i})$ for $i=1,\cdots,n$. Set
$S_{0}:=m(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n})=(mb+\sum_{i=1}^{n}\left\lfloor\frac{ma_{i}}{b_{i}}\right\rfloor)A_{0}+A_{1}+\cdots+A_{n}$
Any other integer $m^{\prime}$ for which the congruences $m^{\prime}a_{i}\equiv 1\,(\text{{\rm mod} }b_{i}),\text{ for }i=1,\cdots,n$ hold can be written $m^{\prime}=m+rl$, where $l$ is the least common multiple of $b_{1},b_{2},\ldots,b_{n}$ and $r\in\mathbb{Z}$. Set $m_{r}=m+rl$. The reader will verify that
(4.4.1) $S_{r}:=m_{r}(bA_{0}+a_{1}A_{1}+\cdots+a_{n}A_{n})=S_{0}+rle(M)A_{0}$
We noted after the statement of Corollary 1.3 that $e(M)=\varepsilon|H_{1}(M)|/b_{1}b_{2}\cdots b_{n}$ for some $\varepsilon\in\{\pm 1\}$. Thus from (4.4.1) we have
(4.4.2) $S_{r}=S_{0}+\varepsilon r\frac{|H_{1}(M)|}{(b_{1}b_{2}\cdots b_{n}/l)}A_{0}$
When $M$ is an integer homology $3$-sphere, $H^{2}(M)=0$, so $\nu_{M}$ is trivial. This is reflected in the fact that $l=b_{1}b_{2}\cdots b_{n}$ in this case, so we can choose $r$ so that the coefficient of $A_{0}$ in $S_{r}$ is 2n2-n.
Similarly, if $H_{1}(M)\cong\mathbb{Z}/2$, the fact that the tangent bundle to the Seifert fibres of $M$ is a trivial line bundle combines with the method of proof of [9, Proposition 2.1] to show that $e(\nu_{M})=0$. We leave it to the reader to verify this using (4.4.2) and Theorem 1.1.

### v2

@@ 3.1. Seifert notation
Consider $M$ constructed by attaching $n+1$ solid tori $N_{i}$, $i=0,1,\dots,n$, to a trivial $S^{1}$-bundle over an orientable surface $F$ with $|\partial F|=n+1$, where we assume that $N_{0}$ is a tubular neighbourhood of a regular fibre and the $N_{i}$ are tubular neighbourhoods of the singular fibres for $i=1,2,\ldots,n$. Denote this trivial $S^{1}$-bundle by $P\cong F\times S^{1}$. Fix orientations on $F$ and on the fibres so that they intersect in $M$.
@@
  3. (3)
    The effect on $b,a_{1},\ldots,a_{n}$ and $e(M)$ of reversing the orientation of $M$ is to replace them by their negatives.
@@
It follows from [13, Theorem 4] that the image of 1-1 under the transgression homomorphism $\delta$ is the Euler class $bu$ of the circle bundle. Hence $\mbox{kernel}(f^{*})$ is generated by $bu$. Then by (3.1.2), $e(\nu_{M})=0$ if and only $e(\tau_{B})=\pm\chi(B)u$ is a multiple of $bu$, which is condition (3.1.1).
@@
    From $\sigma$ one can define an obstruction $2$-cochain, denoted by $c_{\sigma}$, that represents the relative Euler class of the normal bundle $\nu_{M}$ in $H^{2}(M,\sqcup T_{i})$. By construction, the image of $c_{\sigma}]$ under $H^{2}(M,\sqcup T_{i})\to H^{2}(M)$ is the Euler class $e(\nu_{M})$. (See [14, Section 2] for instance.)
@@ 4.1. The second cohomology group of $\Gamma$
Jankins and Neuman determined a canonical isomorphism between $H^{2}(\Gamma)$ and the abelian group having generators $A_{0},A_{1},\ldots,A_{n}$ and relations $A_{0}=b_{i}A_{i}$ for $1\leq i\leq n$ [16]. (The case that $g=0$ is discussed in [16] and the general case in [15].) Each element of $H^{2}(\Gamma)$ can be uniquely expressed in the form
@@ 4.2. The normal bundle of the Seifert fibring of $M$ and discrete representations
$\psi(\gamma)(x,y)=(\gamma\cdot x,\varphi(\gamma)(M))$
@@
Now proceed as in the proof of [3, Lemma A.3], starting from its third paragraph, to show that the $\mathbb{H}^{2}$-bundle $M\times_{\psi}\mathbb{H}^{2}\to M$ is topologically equivalent to the normal bundle $\nu_{M}$ and conclude, as in [3, Proposition A4], that $e(\nu_{M})=e(\varphi)=e(\bar{\varphi}\circ\alpha)$. Thus
@@
It is well-known that $e(\bar{\varphi})$, respectively $e(\varphi)$, is the obstruction to lifting the representation $\bar{\varphi}:\Gamma\to PSL_{2}(\mathbb{R})$, respectively $\varphi:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$, to $\widetilde{PSL}_{2}(\mathbb{R})$ [19, Lemma 2].
@@
We saw in (4.2.1) that that $e(\nu_{M})=\alpha^{*}(e(\bar{\varphi}))=e(\bar{\varphi}\circ\alpha)=e(\varphi)$, where $\varphi=\bar{\varphi}\circ\alpha:\pi_{1}(M)\to PSL_{2}(\mathbb{R})$. Hence, if $e(\nu_{M})=0$, then $\varphi$ lifts to a homomorphism $\pi_{1}(M)\to\widetilde{PSL}_{2}(\mathbb{R})\leq\mbox{Homeo}_{+}(\mathbb{R})$ [19, Lemma 2]. It follows that $\pi_{1}(M)$ is left-orderable [8, Theorem 1.1] and therefore, $M$ admits a horizontal foliation [8, Theorem 1.3]. ∎
5. Seifert fibred examples
@@
That any co-oriented taut foliation on $M$ must be horizontal is [7, Corollary 7] in the case $n=3$ and [1, Proposition 6.1] in general. Thus (1) holds. Consequently, the tangent bundle of any such foliation is isomorphic to $\nu_{M}$ (as an unoriented bundle), which implies (2). ∎
@@
Proof of Theorem 1.7.
Let $K$ be the right-handed trefoil knot, and let $\mu$ and $\lambda$ denote the standard meridian and longitude of $K$. There is a natural Seifert fibre structure on the exterior $X=S^{3}\setminus\nu(K)$, whose regular fibre has slope $6\mu+\lambda$ on $\partial X$. Hence, $p/q$-surgery on $K$ yields a small Seifert fibred space whenever $p/q\neq 6$.
It is known that $K(p/q)$ admits co-orientable taut foliations when $p/q<1$ [25, §3] and hence has a left-orderable fundamental group [4]. On the other hand, assuming that $p>0$ and $p,q$ are coprime, [14, Theorem 1.7] shows that the Euler class of the co-orientable taut foliation constructed by Roberts in [25] vanishes if and only if
@@
This completes the proof. ∎
@@
If $e(\nu_{M})=0$, then $d=\gcd(b_{i},b_{j})$ implies $a_{i}\equiv a_{j}\mbox{ {\rm(mod $d$)}}$.
@@
if they differ by a finite sequence of moves each of which either multiplies the coordinates by 1-1 or is of the form described in Remark 3.1(1). Equivalent sets of Seifert invariants classify Seifert structures on a closed $3$-manifold up to homeomorphism. See Remark 3.1.